# Improving LLM Personas via Rationalization with Psychological Scaffolds

### v1

@@ Abstract
Language models prompted with a user description or persona can predict a user’s preferences and opinions, but existing approaches to building personas—based solely on a user’s demographic attributes and/or prior judgments—fail to capture the underlying reasoning behind said user judgments. We introduce PB&J (Psychology of Behavior and Judgments), a framework that improves LLM personas by incorporating rationales of why a user might make specific judgments. These rationales are LLM-generated, and aim to reason about a user’s behavior on the basis of their experiences, personality traits or beliefs. This is done using psychological scaffolds—structured frameworks grounded in theories such as the Big 5 Personality Traits and Primal World Beliefs—that help provide structure to the generated rationales. Experiments on public opinion and movie preference prediction tasks demonstrate that LLM personas augmented with PB&J rationales consistently outperform methods using only a user’s demographics and/or judgments. Additionally, LLM personas constructed using scaffolds describing user beliefs perform competitively with those using human-written rationales.
Improving LLM Personas via Rationalization
@@ 1 Introduction
Figure 1: Overview of the PB&J framework: A base persona comprising user Demographics and Judgments is augmented with post-hoc rationales of various Scaffolds generated by an LLM. The updated persona is integrated into the system prompt of an LLM to align predictions with user behavior.
@@
Recent advances in prompting large language models (LLMs) to adopt specific user personas offer straightforward yet powerful ways to simulate human behavior and decision-making. A persona is a representation of a single user’s characteristics, preferences, and behavior, designed to guide and align interactions in a personalized system D’Onofrio (2020). Personas have enabled simulation capabilities in LLMs, such as aiding survey design by simulating a wide range of human responses Argyle et al. (2023); Santurkar et al. (2023); Tjuatja et al. (2023), simulating communities to study behavior Park et al. (2022); Zhou et al. (2024); Park et al. (2024), and producing large-scale and diverse synthetic data Moon et al. (2024); Ge et al. (2024). Yet, simulated LLM personas still struggle to align well with intended user behavior Gupta et al. (2024); Liu et al. (2024).
@@
A typical task for evaluating user-LLM persona alignment is opinion prediction. For example, consider opinion-based questions such as “How much are you interested in visiting other countries?”, or “For your job or career aspects of your life, would you say that you are where you expected to be at this point in your life?” A well-aligned LLM persona for a particular user should be able to generate the same responses to these questions as the user. Some approaches to create these personas involve fine-tuning LLMs on user conversations and interaction history Mazaré et al. (2018); Madotto et al. (2019); Li et al. (2024); however, this approach does not scale well due to the lack of adequate user-specific data and the need to fine-tune LLMs for every new user. Zero-shot prompting approaches, where a persona is specified using demographic attributes in a system prompt, avoid fine-tuning but often produce inconsistent responses and expose known biases in current LLMs Santurkar et al. (2023); Hu and Collier (2024); Cheng et al. (2023a); Gupta et al. (2024).
To address these issues, recent approaches incorporate user judgments—prior user responses and history—in the persona prompt Hwang et al. (2023); Sun et al. (2024). For example, a user’s response of “No, not right now” to a question “Are you currently focused on your professional life and career?” is a judgment that could inform their response to the question about where they expected themselves to be career-wise at this moment. While these methods provide additional context, they still fail to address a critical gap: the ability to rationalize why a user might hold a specific judgment, an essential component for understanding behavior. In the context of our example, a rationalization for the user’s answer may be that the user is prioritizing their family over their career at this time. We hypothesize that incorporating post-hoc rationalization for why a user might have made a particular judgment can improve persona prompting in LLMs, bridging the gap between surface-level personas and deeper, more nuanced simulated behavior. However, collecting these rationales is expensive and existing datasets used for persona building are not equipped with them; they only contain user judgments Kirk et al. (2024); Durmus et al. (2024), and in rare cases, user demographics Santurkar et al. (2023); Harper and Konstan. (2015).
In this paper, we test our hypothesis by introducing PB&J (Psychology of Behaviour and Judgements), a framework that uses post-hoc rationales for user judgments, generated by LLMs, to improve LLM personas. PB&J incorporates psychological theories, such as personality traits (e.g., Big 5 traits Goldberg (1993), Schwartz’s value theory Schwartz (1992)) and belief systems (e.g., Primal World Beliefs Clifton et al. (2019)) and uses these theories to guide LLM generated-rationales for given user judgments; we call these psychological scaffolds. It is important to note that these rationales may not reflect the true reasoning behind a user’s judgment, as that is still not present in user information in available datasets. However, we hypothesize that these generated rationales contain plausible markers of real behavioral reasoning due to the presence of such reasoning in the LLM’s training data Binz and Schulz (2023); Hagendorff (2023), as people regularly use everyday reasoning to rationalize behavior — a process known as folk psychology Churchland and Haldane (1988); Malle and Knobe (1997); Malle (2004).
@@
Given a user’s demographics and prior judgments, PB&J uses LLM-generated rationales for user’s judgments to construct a richer, more comprehensive user description for persona prompting. We hypothesize that such richer descriptions help LLMs to move away from making generalizations to more individual responses while predicting user responses Gupta et al. (2024); Hwang et al. (2023). This persona is then used by LLMs to generate predictions which better align with the user’s true responses. We evaluate PB&J on two tasks that represent different aspects of personalization: opinion prediction in OpinionQA Santurkar et al. (2023) and movie preference prediction in MovieLens Harper and Konstan. (2015). Our experiments on both datasets show that LLM personas prompted with PB&J rationales are significantly better at predicting a user’s judgments than those prompted with user demographics and/or prior judgments alone. The improvements from PB&J are seen across nearly all demographic categories, and PB&J can improve personas with or without demographic information. Additionally, we present a pilot experiment where we collect human-written rationales, along with their demographics and judgments on a subset of OpinionQA; LLM-rationales nearly match the performance of human-written rationales in PB&J, showing the efficacy of generated rationales in our approach.
@@ 2 Background
We first introduce the notion of rationalization in the context of reasoning about user behavior, similar to how humans use common explanations to rationale each other’s actions Malle (2004). We hypothesize that rationales are common enough in our discourse to be prevalent in training corpora for LLMs to be able to be used in zero-shot settings. In order to guide LLMs better, we employ various psychological scaffolds, which are templates based on more structured theories of psychology, to rationalize user judgment.
@@ 2.1 Rationalization
Psychologists have long studied how people utilize everyday language to rationalize and anticipate behavioral patterns in social interactions, otherwise known as folk psychology Churchland and Haldane (1988); Malle and Knobe (1997); Malle (2009); Gordon (1986). Building on the example introduced in Section 1, suppose that we have to rationalize why a user believes that they are not focused on their professional life and career at the moment (See Figure 1). We could attribute this behavior to some belief, such as life is dull and uninteresting and chasing career progressions is a waste of time Dennett (1981). Or, we might offer life experiences that shape beliefs, such as personal life issues that are prioritized over career at the moment Malle (2004). For our purposes, we define rationalization as an attempt to understand the behavior of others using explanations based on attributed mental states.
We hypothesize that LLMs have embedded key reasoning required to rationalize user behavior—often present in narrative text (e.g., novels, biographies, social media)—and can thus generate effective rationales Binz and Schulz (2023); Hagendorff (2023). Unlike prior work, which uses rationalization as reasoning chains to explain model decisions post-hoc—either to improve model task performance Wei et al. (2023); Zelikman et al. (2022); Ramnath et al. (2024) or enhance human utility Joshi et al. (2023); Si et al. (2024); Chaleshtori et al. (2024)—we focus on rationalization as a framework for modeling and predicting human behavior.
To rationalize with an LLM, we provide a persona $\mathcal{Q}$ and user’s judgment $j^{i}$ as input and instruct the model to generate “a reasonable explanation that the person would provide for holding that judgment”. The basic rationale $r^{i}_{\mathcal{R}}$ is defined as follows:
$r^{i}_{\mathcal{R}}=LLM_{\mathcal{R}}(j^{i},\mathcal{Q})$ (1)
Here $LLM$ is an off-the-shelf pretrained language model and $\mathcal{R}$ denotes the instructions used to elicit a rationale.
@@ 2.2 Psychological scaffolds
While Equation 1 helps us generate rationales from LLMs, it is important to control what LLMs generate as plausible reasons for human behavior. We believe that LLMs can provide structured rationales that incorporate well-established psychological theories. We call these psychological scaffolds as they extend the LLMs’ reasoning power and reach by providing coherent, theoretical scaffolds for reasoning about observed behavior.
@@
The specific scaffolds we study in this work are shown in Table 1. In the context of our previous example, the experiences scaffold might promote a rationale based on the bad experience of a family member who prioritized career goals over personal life McAdams (1993). Rationales generated from the personality traits scaffold may posit the user’s openness to new experiences, including bold career choices Goldberg (1993); Schwartz (1992). A rationale using the Primal World Beliefs scaffold may surmise that the user sees the world as dull and uninteresting and thus not worth the effort of seeking professional success Clifton et al. (2019). More details about the theories underlying each scaffold can be found in Appendix H. These scaffolds offer diverse yet complementary ways to structure psychological rationales, anchoring explanations in personal experiences, personality traits, and belief systems. By providing a systematic way to interpret behavior, we show that scaffolds enhance the coherence and depth of rationalization.
@@
To generate a rationale from an LLM with a particular scaffold, we include additional scaffold information in the rationale generation instructions to arrive at scaffold-specific instructions $\psi$. Using $\mathcal{Q}$ and $j^{i}$ as above, the scaffold rationale $r^{i}_{\psi}$ is defined as follows:
$r^{i}_{\psi}=LLM_{\psi}(j^{i},\mathcal{Q})$ (2)
The specific instructions used for each scaffold are given in Appendix A.
3 LLM Personas using PB&J
@@
Consider a task where an LLM has to adopt the persona of a given user, and perform a task (e.g. answer opinion-based questions) as their proxy. In this section, we describe how we leverage LLM rationalization to construct LLM personas, which are used as prompts to generate personalized responses as a proxy for the user. We begin by setting up a basic persona description composed of user demographics and augmented with a set of user judgments as in previous work Hwang et al. (2023). We then present PB&J, which enriches persona descriptions by providing additional context in the form of rationales, and demonstrate how PB&J is used to generate personalized responses as a proxy for the user.
3.1 Persona Building
Base Persona Setup.
Following Hwang et al. (2023), we begin with a basic persona $\mathcal{Q}_{B}$ for an arbitrary user consisting of demographic attributes $\mathcal{D}$ and a set of $N$ seed judgments $\mathcal{J}$. Demographic attributes $\mathcal{D}$ capture the sociological characteristics and group identity traits of a user, such as age, gender, education, and race Santurkar et al. (2023). While these attributes can provide useful information on user preferences and opinions, prior work has shown that relying solely on demographics often results in stereotypical and biased responses from LLMs Hwang et al. (2023); Cheng et al. (2023b); Gupta et al. (2024). To address these limitations, Hwang et al. (2023) augment demographic attributes with a set of seed judgments $\mathcal{J}$. These seed judgments often represent user-selected answers from a held out set of questions, or a prior interaction history which aims to provide additional context to enrich the persona description Hwang et al. (2023); Do et al. (2024). As illustrated in Figure 1, the user’s response to the question “Are you currently focused on your professional life and career?” is a judgment. In other tasks, a user’s interaction history, such as the movies they watched beforehand along with their reviews, would be seed judgments.
Post-hoc Rationales for Seed Judgments.
In PB&J, we augment the above persona description $\mathcal{Q}_{B}$ with additional context in the form of post-hoc generated rationales $r^{i}$ for each seed judgment $j^{i}\in\mathcal{J}$ following Equation 1. Each post-hoc rationale is generated independently for the corresponding seed judgment and is added to the persona alongside the judgment itself. Our rationalized persona is defined as:
$\mathcal{Q}_{\mathcal{R}}=(\mathcal{D};\{(j^{i},r^{i}_{\mathcal{R}}):0\leq i\leq N\}).$ (3)
Building on Section 2.1, post-hoc rationales operationalize how $LLM_{\mathcal{R}}$ generates psychological rationales to justify why a user might hold a given judgment. Assuming that LLMs are trained on a sufficiently large corpora of human-written text—likely containing a surfeit of human rationales about behavior—they should be able to generate similar language constructs to rationalize user behavior without requiring anthropomorphism. [^1] This does not imply that LLMs actually understand user behavior; it is possible that LLMs mimic reasoning patterns found in their training data.
If we use Equation 2 to generate rationales instead, we arrive at the scaffold rationalized persona:
$\mathcal{Q}_{\mathcal{\psi}}=(\mathcal{D};\{(j^{i},r^{i}_{\mathcal{\psi}}):0\leq i\leq N\}).$ (4)
We observe that strong language models tend to follow the rationalization instructions provided in $\psi$.
3.2 Generating responses from Persona LLMs
Similar to prior work Santurkar et al. (2023); Hwang et al. (2023), persona descriptions $\mathcal{Q}$ obtained in Section 3.1 can be used as prompts to steer responses towards a given user persona. Consider a task where an individual user has provided several input-output pairs $\{x^{i},y^{i}\}$ (e.g., opinion question and answer pairs). The persona description can provide additional context to these inputs, enabling an LLM to generate predictions that align with the user’s preferences and behavior.
We predict a user’s output for an input $x^{i}$ by providing task-specific instructions $\mathcal{T}$ to an off-the-shelf language model along with the user’s persona $Q$.
$\hat{y}^{i}=LLM_{\mathcal{T}}(x^{i},\mathcal{Q}),$ (5)
To evaluate a persona building method, we vary $\mathcal{Q}$ over the base, rationalized, and scaffold rationalized personas $\mathcal{Q}_{B},\mathcal{Q}_{\mathcal{R}},\mathcal{Q}_{\psi}$. We consider aggregate prediction accuracy under this model across test users as a measure of the fidelity of the persona building method.
4 Experiments
@@ C. Refused to answer
Task and Datasets.
We conduct experiments on two datasets - OpinionQA Santurkar et al. (2023) and MovieLens Harper and Konstan. (2015). OpinionQA is a collection of American public opinion surveys conducted by PEW Research containing user-selected answers to multiple-choice questions in 15 different topics ranging from food safety to guns. The dataset also contains multiple demographic attributes for every user. Our task is to predict a user’s answer for a question, given their persona description. For OpinionQA, we use a subset of 750 users along with 10 test questions for each user. MovieLens is a movie rating dataset containing timestamped movie ratings corresponding to individual users. While the demographic information in MovieLens is not as rich as OpinionQA, we use an older subset of the dataset that records the gender, age, occupation and location of each user. Our task is to predict a user’s rating, between 1-5, for a movie, given their persona description. For MovieLens, we use a subset of 100 users, with 10 test movies for each user. Further details and splits of the dataset are presented in Appendix B. Example inputs-outputs of each of these datasets are shown in Figure 2.

### v2

# Improving Language Model Personas via Rationalization with Psychological Scaffolds

@@ Abstract
Language models prompted with a user description or persona are being used to predict the user’s preferences and opinions. However, existing approaches to building personas mostly rely on a user’s demographic attributes and/or prior judgments, but not on any underlying reasoning behind a user’s judgments. We introduce PB&J (Psychology of Behavior and Judgments), a framework that improves LM personas by incorporating potential rationales for why the user could have made a certain judgment. Our rationales are generated by a language model to explicitly reason about a user’s behavior on the basis of their experiences, personality traits, or beliefs. Our method employs psychological scaffolds: structured frameworks such as the Big 5 Personality Traits or Primal World Beliefs to help ground the generated rationales in existing theories. Experiments on public opinion and movie preference prediction tasks demonstrate that language model personas augmented with PB&J rationales consistently outperform personas conditioned only on user demographics and / or judgments, including those that use a model’s default chain-of-thought, which is not grounded in psychological theories. Additionally, our PB&J personas perform competitively with those using human-written rationales, suggesting the potential of synthetic rationales guided by existing theories.
Improving Language Model Personas via Rationalization
@@ 1 Introduction
Figure 1: Overview of the PB&J framework: A base persona comprising user Demographics and Judgments is augmented with post-hoc rationales of various Scaffolds generated by an LM. The updated persona is integrated into the system prompt of an LM to align predictions with user behavior.
@@
Recent advances in language modeling for user persona (i.e, characteristics, preferences, and behavior of a user) offer straightforward yet powerful ways to predict user behavior and decision-making D’Onofrio (2020). Personas have enabled simulation capabilities in LMs, such as helping survey design by simulating a wide range of human responses Argyle et al. (2023); Santurkar et al. (2023); Tjuatja et al. (2023), simulating communities to study behavior Park et al. (2022); Zhou et al. (2024); Park et al. (2024), and producing diverse and large-scale synthetic data Moon et al. (2024); Ge et al. (2024). Yet, simulated LM personas still struggle to align well with intended user behavior Gupta et al. (2024); Liu et al. (2024).
@@
A typical task for evaluating LM personas is opinion prediction: how accurately does the LM persona reflect the user’s real opinion to questions such as “For your job or career aspects of your life, would you say that you are where you expected to be at this point in your life?” Fine-tuning LMs on user conversations and interaction history to create personas does not scale well due to the lack of adequate user-specific data Mazaré et al. (2018); Madotto et al. (2019); Li et al. (2024). Zero-shot prompting approaches, where a persona is specified using demographic attributes in a system prompt, avoid these issues but often produce inconsistent responses and expose known biases in current LMs Santurkar et al. (2023); Hu and Collier (2024); Cheng et al. (2023a); Gupta et al. (2024). To improve zero-shot prompting, recent approaches incorporate user judgments (prior user responses and history) in the prompt Hwang et al. (2023); Sun et al. (2024). For example, a user judgment like “I’m not very focused on my professional life right now” could inform their response to questions about their career expectations. While these methods provide additional context, they still fail to address a critical gap: the ability to rationalize why a user might have a specific judgment, an essential component to understand behavior. In our example, one plausible rationale for the user’s judgment could be that they are prioritizing their family over their career at this time.
In this work, we hypothesize that incorporating plausible post-hoc rationales for a user’s judgment could improve LM personalization, bridging the gap between surface-level personas and deeper, more nuanced simulated behavior. However, collecting such rationales is expensive and existing datasets for building personas only contain judgments Kirk et al. (2024); Durmus et al. (2024), and in rare cases, user demographics Santurkar et al. (2023); Harper and Konstan. (2015), but no rationales for judgments. To address this limitation, we introduce PB&J (Psychology of Behavior and Judgments), a framework that uses post-hoc plausible, yet synthetic, LM-generated rationales for user judgments, to improve LM personas. The rationales for user judgments in our method stem from psychological theories, such as personality traits (e.g., Big 5 traits Goldberg (1993), Schwartz’s value theory Schwartz (1992)) and belief systems (e.g., Primal World Beliefs Clifton et al. (2019)); we call these psychological scaffolds. Each psychological scaffold offers additional context to produce one potential (albeit generated) path of reasoning to support a given user judgment. This is in contrast to chain-of-thought rationales Wei et al. (2023) where no explicit psychological scaffold is employed; the LM relies on its own world knowledge to rationalize a test input prediction ad hoc.
It is important to note that while our synthetic rationales may not reflect the true reasoning behind a user’s judgment, they may contain plausible markers of real behavioral reasoning due to the presence of such reasoning in the LM’s training data Binz and Schulz (2023); Hagendorff (2023); Balepur et al. (2025). Indeed, humans use everyday reasoning to rationalize behavior; this is known as folk psychology Churchland and Haldane (1988); Malle and Knobe (1997); Malle (2004).
@@
Given a user’s demographics and prior judgments, PB&J produces LM-generated rationales for user’s judgments to construct a richer, more comprehensive user description for persona prompting. Our experiments show that when prompted with such richer descriptions, LMs can result in more accurate personas, as evaluated on a set of test instances. We evaluate PB&J on two tasks that represent different aspects of personalization: opinion prediction in OpinionQA Santurkar et al. (2023) and movie preference prediction in MovieLens Harper and Konstan. (2015). On both benchmarks, we show that LM personas prompted with PB&J rationales are significantly better at predicting a user’s response to test instances than those prompted with user demographics and/or prior judgments alone, including with additional chain-of-thought reasoning. Improvements from PB&J are seen across nearly all demographic categories, and PB&J can improve personas with or without demographic information. Finally, we present a pilot experiment where we collect human-written rationales, along with their demographics and judgments on a subset of OpinionQA; LM-rationales nearly match the performance of human-written rationales in PB&J, showing the efficacy of generated rationales in our approach.
At a high level, our work highlights the value of synthesizing data with the appropriate structural context to produce plausible rationalization that could improve the modeling of user opinions and preferences in already capable language models.
@@ 2 Background
Humans use common explanations to rationale each other’s actions Malle (2004). We hypothesize that rationales are common enough in our discourse (e.g., narrative text of novels, biographies, social media) to be prevalent in training corpora for LMs to generate effective rationales in zero-shot settings Binz and Schulz (2023); Hagendorff (2023). In order to better guide LMs, we employ various psychological scaffolds, which are templates based on more structured theories of psychology, to rationalize user judgment.
@@ 2.1 Rationalization
Psychologists have long studied how people use everyday language to rationalize and anticipate behavioral patterns in social interactions, also known as folk psychology Churchland and Haldane (1988); Malle and Knobe (1997); Malle (2009); Gordon (1986). Building on the example introduced in Section 1, suppose that we have to rationalize why a user believes that they are not focused on their professional life and career at the moment (See Figure 1). We could assume that this behavior arises due to some belief, such as life is dull and uninteresting and chasing career progressions is a waste of time Dennett (1981). Or we may offer life experiences that shape beliefs, such as personal life issues that get prioritized over one’s career at the moment Malle (2004). For our purposes, we define rationalization as an attempt to understand the behavior of others using explanations based on assumed mental states.
Unlike prior work, which uses rationalization as reasoning chains to explain model decisions ad hoc, to improve model task performance Wei et al. (2023); Zelikman et al. (2022); Ramnath et al. (2024), or to enhance human utility Joshi et al. (2023); Si et al. (2024); Chaleshtori et al. (2024), we focus on post-hoc rationalization of user judgments to improve LM personas for accurate user predictions and preferences. Formally, given a user persona $\mathcal{Q}$ and judgment $j$ as input, we instruct an LM to generate “a reasonable explanation that the user would provide for holding that judgment.” The basic rationale $r_{\text{basic}}$ is defined as
$r_{\text{basic}}=LM_{\mathcal{R}}(j,\mathcal{Q}),$ (1)
where $LM$ is an off-the-shelf pretrained language model instructed to elicit a rationale.
@@ 2.2 Psychological scaffolds
While Equation 1 helps us generate rationales from LMs, it is important to control what LMs generate as plausible rationales for human behavior. Can LMs generate more informed rationales based on well-established psychological theories? To find out, we employ psychological scaffolds to enhance the LMs’ reasoning capabilities about user beliefs and preferences by providing coherent scaffolds as instructions to follow for reasoning about the underlying causes of observed behavior.
@@
The specific scaffolds we study in this work are shown in Table 1. In the context of our previous example, the experiences scaffold might promote a rationale based on the bad experience of a family member who prioritized career goals over personal life McAdams (1993). Rationales generated from the personality traits scaffold may posit the user’s openness to new experiences, including bold career choices Goldberg (1993); Schwartz (1992). A rationale using the Primal World Beliefs scaffold may surmise that the user sees the world as dull and uninteresting and thus not worth the effort of seeking professional success Clifton et al. (2019). More details about the theories underlying each scaffold can be found in Appendix A. While the scaffolds that we select in this work are common psychological constructs used in prior work on personality and social intelligence Zhou et al. (2024); Vu et al. (2022b); Moon et al. (2024), any scaffold that helps reason about a user’s belief can be used. Our selected scaffolds offer diverse yet complementary ways to structure psychological rationales, anchoring rationales in personal experiences, personality traits, and belief systems.
@@
To generate a rationale from an LM with a particular scaffold, we include additional scaffold information in the rationale generation instructions to arrive at scaffold-specific instructions $\psi$. Using $\mathcal{Q}$ and $j$ as above, the scaffold rationale $r_{\psi}$ is defined as follows:
$r_{\psi}=LM_{\mathcal{R}}(j,\mathcal{Q},\psi)$ (2)
The specific instructions used for each scaffold are given in Appendix B.
3 LM Personas using PB&J
@@
Consider a task where an LM has to adopt the persona of a given user, and perform a task (e.g. answer opinion-based questions) as their proxy. In this section, we describe how we leverage LM rationalization to construct LM personas, which are used as prompts to generate personalized responses as a proxy for the user. We begin by setting up a basic persona description composed of user demographics and augmented with a set of user judgments, just as in previous work Hwang et al. (2023). We then present PB&J, which enriches persona descriptions by providing additional context in the form of rationales, and demonstrate how PB&J is used to generate personalized responses as a proxy for the user.
@@ C. Refused to answer
3.1 Persona Building
Base Persona Setup.
Following Hwang et al. (2023), we begin with a basic persona $\mathcal{Q}_{B}$ for an arbitrary user consisting of demographic attributes $\mathcal{D}$ and a set of $N$ seed judgments $\mathcal{J}$. Demographic attributes $\mathcal{D}$ capture the sociological characteristics and group identity traits of a user, such as age, gender, education, and race Santurkar et al. (2023). While these attributes can provide useful information on user preferences and opinions, prior work has shown that relying solely on demographics often results in stereotypical and biased responses from LMs Hwang et al. (2023); Cheng et al. (2023b); Gupta et al. (2024). To address these limitations, Hwang et al. (2023) augment demographic attributes with a set of seed judgments $\mathcal{J}$. Seed judgments often represent user-selected answers from a held out set of questions, or a prior interaction history which aims to provide additional context to enrich the persona description Hwang et al. (2023); Do et al. (2024). In other tasks, a user’s interaction history, such as the movies they watched beforehand along with their reviews, constitute the seed judgments.
Post-hoc Rationales for Seed Judgments.
In PB&J, we augment the description of the above persona $\mathcal{Q}_{B}$ with additional context in the form of post-hoc generated rationales $r$ for each seed judgment $j\in\mathcal{J}$ following Equation 1. Each post-hoc rationale is generated independently for the corresponding seed judgment and is added to the persona alongside the judgment itself. Our rationalized persona is defined as:
$\mathcal{Q}_{\mathcal{R}}=(\mathcal{D};\{(j,r_{\text{basic}}):j\in\mathcal{J}).$ (3)
Building on Section 2.1, post-hoc rationales operationalize how $LM_{\mathcal{R}}$ generates rationales to explain why a user might hold a given judgment.
Using rationales from Equation 2, we arrive at the scaffold rationalized persona,
$\mathcal{Q}_{\mathcal{\psi}}=(\mathcal{D};\{(j,r_{\mathcal{\psi}}):j\in\mathcal{J}).$ (4)
3.2 Generating responses from Persona LMs
Persona descriptions $\mathcal{Q}$ obtained in §3.1 can be used as prompts to steer responses towards a given user persona, following prior work Santurkar et al. (2023); Hwang et al. (2023). Consider a task where an individual user has provided several input-output pairs for a test instance $\{x,y\}$ (e.g., opinion question and answer pairs). The persona description can provide additional context to predict responses for inputs, enabling an LM to generate accurate predictions for a user.
We predict a user’s output for an input $x$ by prompting an off-the-shelf language model $LM_{\mathcal{T}}$ with task-specific instructions $\mathcal{T}$.
$\hat{y}=LM_{\mathcal{T}}(x,\mathcal{Q}),$ (5)
To evaluate a persona building method, we vary $\mathcal{Q}$ over the base, rationalized, and scaffold rationalized personas $\mathcal{Q}_{B},\mathcal{Q}_{\mathcal{R}},\mathcal{Q}_{\psi}$.
4 Experiments